# High CO/H2 ratios supports an exocometary origin for a CO-rich debris disk

### v2

@@ Abstract
Methods. We present near-IR CRIRES+ spectra of HD 110058 and HD 131488 which provide the first direct probe of H2 gas in CO-rich exocometary belts. We target the H2 (v=1-0 S(0)) line at 2223.3 nm and and the 12CO $v=2\rightarrow 0$ rovibrational lines in the range 2333.8-2335.5 nm and derive constraints on column densities along the line-of-sight to the stars.
Results. We strongly detect 12CO but not H2 in the CRIRES+ spectra. This allows us to place $3\sigma$ lower limits on the $\frac{\text{CO}}{\text{H}_{2}}$ ratios of $1.35\times 10^{-3}$ and $3.09\times 10^{-5}$ for HD 110058 and HD 131488 respectively. These constraints demonstrate that at least for HD 110058, the exocometary gas is compositionally distinct and significantly H2-poor, compared to the $10^{-4}$ $\frac{\text{CO}}{\text{H}_{2}}$ ratios typical of protoplanetary disks. We also find H2 alone is unlikely to shield CO over the lifetime of the systems. Overall this suggests that the gas in CO-rich belts is most likely not primordial in origin, supporting the presence of exocometary gas.
@@ 1 Introduction
Exocometary belts, also know as debris disks, are circumstellar rings composed mainly of dust and rocky/icy planetesimals which have optically thin continuum emission, in contrast to protoplanetary disks (Pearce, 2024; Wyatt, 2020; Hughes et al., 2018, provide reviews of the field). Exocometary belts are typically detected through the thermal infrared excess they produce over the stellar photosphere. They can be imaged at shorter wavelengths through scattered light from small dust grains with instruments like the Gemini Planet Imager/Sphere (Crotts et al., 2024; Desgrange et al., 2025) and in mm emission with interferometers like The Atacama Large Millimeter/submillimeter Array (ALMA) (Matrà et al., 2025). Dust in exocometary belts is continually removed by radiation pressure, stellar winds, and/or Poynting-Robinson drag (Backman & Paresce, 1993) and replenished by destructive collisional processes between larger grains and planetesimals (Dohnanyi, 1969). Thus exocometary dust is thought to be secondary in origin, rather than a remnant of the belt’s progenitor protoplanetary disk.
@@
Exocometary belts were initially assumed to be gas-free; however a variety of atomic species (eg. Hobbs et al., 1985) as well as CO have been detected in over 20 exocometary systems (Brennan et al., 2025; Cataldi et al., 2023; Marino et al., 2020; Kral et al., 2017; Moór et al., 2017; Matrà et al., 2017b; Lieman-Sifry et al., 2016; Kóspál et al., 2013; Troutman et al., 2011; Dent et al., 2005; Zuckerman, 1995, eg.). Thus far CO is the only molecule detected in exocometary gas due to its resistance to photodissociation, chemical stability, and easy detectability (Visser et al., 2009). CO has strong rovibronic transitions, and the typically cold gas present in exocometary belts is detectable in UV rovibronic absorption spectra by the Hubble Space Telescope (HST) (eg. Roberge et al., 2000), rotational transitions in emission detectable in mm with ALMA (eg. Matrà et al., 2017b; Cataldi et al., 2023; Rebollido et al., 2022) and IR absorption from rovibrational transitions as detected using the Gemini South Telescope and the NASA Infrared Telescope Facility around $\beta$ Pictoris (Troutman et al., 2011).
@@
CO emission from rotational transitions observed with ALMA has shown exocometary belts can be separated into 2 categories of: CO-rich ($M_{\mathrm{CO}}\gtrsim 10^{-3}M_{\bigoplus}$) and CO-poor ($M_{\mathrm{CO}}\lesssim 10^{-3}M_{\bigoplus}$) (Marino et al., 2020). This bimodal distribution of exocometary CO masses can be seen in Cataldi et al. (2023), with CO masses measured directly through 12C16O lines (if optically thin) or through 13C16O or 12C18O measurements (if optically thick and assuming interstellar isotopologue ratios). This gas could be primordial if persisting from the parent protoplanetary disk (Nakatani et al., 2023; Kóspál et al., 2013) or secondary if produced in-situ by outgassing from exocomets (Zuckerman & Song, 2012). Emission from OI has also been detected in the far IR by (Riviere-Marichalar et al., 2014).
@@
When determining the origin of the CO observed in exocometary belts, it is important to consider the role of UV shielding, which attenuates UV photons and lengthens the time taken for CO to photodissociate (typical photodissociation timescales can be found in Visser et al., 2009; Heays et al., 2017). Unlike in protoplanetary disks, the dust in exocometary belts is optically thin at all wavelengths (e.g Matrà et al., 2018a). Thus, it cannot provide circumstellar gas with substantial shielding from the UV radiation due to the host star and interstellar radiation field (ISRF).
@@
If the gas is primordial, the abundant H2 inherited from the protoplanetary disk could provide shielding (Kóspál et al., 2013). If the gas is secondary, CO can be replenished by outgassing from exocomets in the belt (Zuckerman & Song, 2012), and shielding could be provided by CO itself (self-shielding) and/or atomic carbon (Matrà et al., 2017a; Kral et al., 2019), which could be produced by photodissociation of carbon-bearing parent molecules such as CO but also yet unseen species such as CO2 and CH4. In CO-poor belts, the main way to distinguish between the primordial and secondary scenarios has so far been to compare the CO photodissociation timescale to the system age, with longer timescales allowing for a primordial origin and shorter implying secondary gas. For CO-rich belts, the origin scenario is unclear but a key signature of primordial gas would be a H2 abundance comparable with protoplanetary disks.
@@
UV shielding of CO by self-shielding, CI, and H2 has traditionally been thought to be important to the longevity of CO in circumstellar disks. In CO-rich debris disks, it is possible that primordial H2 abundance, assumed through a protoplanetary like $\frac{\text{CO}}{\text{H}_{2}}$ ratio could shield the gas for a few Myr. However, for CO-poor belts numerous studies show that even assuming generously low $\frac{\text{CO}}{\text{H}_{2}}$ ratios of $10^{-6}$, provide insufficient shielding for primordial CO to persists to the typically older ages of these systems (e.g Marino et al., 2016; Matrà et al., 2017b; Marino et al., 2018; Matrà et al., 2019; Kral et al., 2019). This implies that gas in CO-poor systems, from younger systems like $\beta$ Pictoris (Matrà et al., 2017a) to older systems like Fomalhaut (Matrà et al., 2017b) and $\eta$ Corvi (Marino et al., 2016), is most likely secondary. This was corroborated by upper limits on the H2 density in the $\beta$ Pictoris disk obtained from non-local thermodynamic equilibrium (non-LTE) modelling of the CO excitation (Matrà et al., 2017a).
@@
In CO-rich belts, on the other hand, the origin of the gas remains an open question. These disks have CO masses comparable with the lower end observed in Herbig-Ae protoplanetary disks around stars that are several Myr to a few 10s of Myr old (Moór et al., 2020). If the gas in these CO-rich disks is primordial, we would expect a large amount of H2, the most abundant molecular species in primordial gas. In protoplanetary disks, the gas phase CO-to-H2 ratio appears to decrease over time, producing $\frac{\text{CO}}{\text{H}_{2}}$ ratios between $10^{-4}$ and $10^{-6}$ (see Figure 5, Table 4 and Bergin & Williams (2018); Zhang et al. (2020)). This is because CO abundances start close to ISM-like, but overtime CO can be processed into other carbon-carriers (such as CO2 and CH3OH) and at the same time CO can be sequestered onto icy grains as they grow and settle to the mid-plane (Schwarz et al., 2018; Bosman et al., 2018; Krijt et al., 2018, 2020). For comparison, in Solar System comets, $\frac{\text{CO}}{\text{H}_{2}}$ ratios as high as 0.73 have been observed as H2 is mainly produced as a product of H2O photodissociation (Bockelée-Morvan & Biver, 2017; Feldman et al., 2002). Thus, in exocometary belts, a $\frac{\text{CO}}{\text{H}_{2}}$ abundance ratio higher than 10-4 is diagnostic of a secondary gas origin and this ratio can be used to calculate whether there is sufficient H2 shielding to allow persistence of CO to the ages of these systems without replenishment or other shielding mechanisms.
@@
Some constraints on the presence of H2 in exocometary belts are present in the literature. For example, using HST Lecavelier des Etangs et al. (2001) constrained upper limits for the column density of H2 along the line of sight to the edge-on, CO-poor $\beta$ Pictoris disk to be ¡ 0.1 M and combined this with a CO column density (Roberge et al., 2000) to find a $\frac{\text{CO}}{\text{H}_{2}}$ ratio lower limit of $6\times 10^{-4}$, higher than expected of primordial gas. Their conclusion is in agreement with the conclusions of Matrà et al. (2017a) that a primordial origin scenario for the gas in $\beta$ Pictoris is unlikely as there is insufficient H2 present to provide shielding for the lifetime of the system. H2 can also be indirectly constrained by gas kinematics/distribution. Fernandez et al. (2005) uses the breaking experienced by ionised metallic gas due to unseen neutral H2 gas to estimate the H2 mass in $\beta$ Pictoris. Hughes et al. (2017) uses the scale height of the gas in 49 Ceti to infer an upper limit for H2, via the mean molecular weight of the gas required by models to reproduce the observed scale height for CO.
@@
Hot H2 emission has also been detected in systems bearing debris disks such as TWA 7 (Flagg et al., 2021) and AU Mic (Flagg et al., 2022), where the latter has no evidence for CO (Cronin-Coltsmann et al., 2023). It is unclear whether this hot H2 could originate from star spots or an inner disk of gas accreting onto the star, unrelated to the outer disks. Other works have attempted to detect warm $\text{H}_{2}$ with Spitzer, but these observations lacked the SNR/sensitivity to confirm or refute a primordial origin scenario (Chen et al., 2007; Kóspál et al., 2013).
We use the newly upgraded, high-resolution CRIRES+ instrument on the VLT to observe the edge-on, CO-rich disks around HD 110058 and HD 131488. With a spectral resolution $R\sim 100000$ (Leibundgut et al., 2022), CRIRES+ is ideal for detecting line of sight absorption due to CO and H2. Our observations cover multiple rovibrational lines of 12CO from the v=2-0 level. Absorption from CO in edge-on debris disks has been successfully observed by Troutman et al. (2011), who reported low-$J$ CO absorption from the $\beta$ Pictoris disk using IRTF/CSHELL. Worthen et al. (2024) reported a non-detection of CO in absorption toward the edge-on HD 32297 disk using IRTF/iSHELL which was used to estimate the scale height of the disk and place an upper limit on the line-of-sight CO column density. The two disks in this study were measured in absorption in the UV using Hubble by Brennan et al. (2025). Our CRIRES+ observations allow for CO column density and temperature estimates for both HD 110058 and HD 131488.
@@
In this work, we aim to directly probe the column density of H2 within CO-rich exocometary belts for the first time, and in doing so establish the CO origin. We expect any exocometary $\text{H}_{2}$ to be co-located with and similar temperature as the CO detected through UV absorption spectroscopy in both disks by Brennan et al. (2025) and in IR in this work. However it is worth noting that H2 should have a larger vertical extent than CO due to CO being more readily photodissociated at the disk’s surface and it’s lower mean molecular weight (Marino et al., 2020; Hughes et al., 2017). Both H2 and CO should be cold and produce narrow absorption lines from their ground vibrational energy levels. We focus on absorption spectroscopy in edge-on belts to use the star as a bright background continuum, and maximise the column density of gas along the line of sight to it. We identified the H2 v=1-0 S(0) line at 2223 nm as the most promising transition probing rovibrational absorption from the ground level of the H2 molecule. This is because, assuming the H2 gas in these systems has a similar temperature to the CO and is in LTE, this line would be the strongest IR transition.
@@
Stellar and disk properties for HD 110058 and HD 131488 are detailed in Table 1. Both are A-type stars (the most common spectral type for CO-rich belts, Moór et al., 2020) and thus provide bright stellar continua against which we searched for absorption. The disks in these systems are also very close to edge-on with inclinations of $85.5^{+2.5}_{-7.2}$ and $82\pm 3$ respectively (Moór et al., 2017; Hales et al., 2022). We choose these two young ($\sim 15$ Myr), CO-rich, edge-on belts as they have the highest CO column densities detected in absorption to date (Brennan et al., 2025).

### v3

# High CO/H2 ratio supports an exocometary origin for a CO-rich debris disk

@@ Abstract
Methods. We present near-IR CRIRES+ spectra of HD 110058 and HD 131488 which provide the first direct probe of H2 gas in CO-rich exocometary belts. We target the H2 (v=1-0 S(0)) line at 2223.3 nm and the 12CO $v=2\rightarrow 0$ rovibrational lines in the range 2333.8-2335.5 nm and derive constraints on column densities along the line-of-sight to the stars.
Results. We strongly detect 12CO but not H2 in the CRIRES+ spectra. This allows us to place $3\sigma$ lower limits on the $\frac{\text{CO}}{\text{H}_{2}}$ ratios of $1.35\times 10^{-3}$ and $3.09\times 10^{-5}$ for HD 110058 and HD 131488 respectively. These constraints demonstrate that at least for HD 110058, the exocometary gas is compositionally distinct and significantly H2-poor, compared to the $10^{-4}$ $\frac{\text{CO}}{\text{H}_{2}}$ ratios typical of protoplanetary disks. For HD 131488, we further compare the CO photodissociation timescale to the age of the system through simple shielding arguments, finding that we cannot formally rule out a primordial origin; however, we suggest that a more realistic model of CO survival is likely to support a secondary origin for this system as well. Overall, a high $\frac{\text{CO}}{\text{H}_{2}}$ ratio for HD 110058 indicates that the gas in this CO-rich belt is most likely not primordial in composition, supporting the presence of exocometary gas.
@@ 1 Introduction
Exocometary belts, also known as debris disks, are circumstellar rings composed mainly of dust and rocky/icy planetesimals which have optically thin continuum emission, in contrast to protoplanetary disks (Pearce 2026; Wyatt 2020; Hughes et al. 2018, provide reviews of the field). Exocometary belts are typically detected through the thermal infrared excess they produce over the stellar photosphere. They can be imaged at shorter wavelengths through scattered light from small dust grains with instruments such as the Gemini Planet Imager/Sphere (Crotts et al. 2024; Desgrange et al. 2025) and in mm emission with interferometers such as The Atacama Large Millimeter/submillimeter Array (ALMA) (Matrà et al. 2025). Dust in exocometary belts is continually removed by radiation pressure, stellar winds, and/or Poynting-Robinson drag (e.g. Strubbe and Chiang 2006; Backman and Paresce 1993) and replenished by destructive collisional processes between larger grains and planetesimals (Dohnanyi 1969). Thus, exocometary dust is thought to be secondary in origin, rather than a remnant of the belt’s progenitor protoplanetary disk.
@@
Exocometary belts were initially assumed to be gas-free; however a variety of atomic species (e.g. Hobbs et al. 1985) as well as CO have been detected in over 20 exocometary systems (e.g. Brennan et al. 2024; Cataldi et al. 2023; Marino et al. 2020; Kral et al. 2017; Moór et al. 2017; Matrà et al. 2017b; Lieman-Sifry et al. 2016; Kóspál et al. 2013; Troutman et al. 2011; Dent et al. 2005; Zuckerman et al. 1995). Thus far CO is the only molecule detected in exocometary gas due to its resistance to photodissociation, chemical stability, and easy detectability (Visser et al. 2009). CO has strong rovibronic transitions, and the typically cold gas present in exocometary belts is detectable in UV rovibronic absorption spectra by the Hubble Space Telescope (HST) (e.g. Roberge et al. 2000), rotational transitions in emission detectable in mm with ALMA (e.g. Matrà et al. 2017b; Cataldi et al. 2023; Rebollido et al. 2022) and IR absorption from rovibrational transitions as detected using the Gemini South Telescope and the NASA Infrared Telescope Facility around $\beta$ Pictoris (Troutman et al. 2011).
@@
CO emission from rotational transitions observed with ALMA has shown that exocometary belts can be separated into 2 categories of: CO-rich ($M_{\mathrm{CO}}\gtrsim 10^{-3}M_{\bigoplus}$) and CO-poor ($M_{\mathrm{CO}}\lesssim 10^{-3}M_{\bigoplus}$) (Marino et al. 2020). This bimodal distribution of exocometary CO masses can be seen in Cataldi et al. (2023), with CO masses measured directly through 12C16O lines (if optically thin) or through 13C16O or 12C18O measurements (if optically thick and assuming interstellar isotopologue ratios). This gas could be primordial if persisting from the parent protoplanetary disk (Nakatani et al. 2023; Kóspál et al. 2013) or secondary if produced in-situ by outgassing from exocomets (Zuckerman and Song 2012).
@@
When determining the origin of the CO observed in exocometary belts, it is important to consider the role of UV shielding of CO by self-shielding, CI, and H2, which attenuates UV photons and lengthens the time taken for CO to photodissociate (Heays et al. 2017; Visser et al. 2009). Unlike in protoplanetary disks, the dust in exocometary belts is optically thin at all wavelengths (e.g Matrà et al. 2018a). Thus, it cannot provide circumstellar gas with substantial shielding from UV radiation due to the host star and interstellar radiation field (ISRF) with the expected lifetime of an unshielded CO molecule in the ISRF being $\sim$ 130 years (Heays et al. 2017). If the gas in an exocometary belt has a photodissociation time shorter than the age of the system, this implies the gas is secondary, as it likely is being replenished by comets in the belt. Furthermore, CO-rich belts require shielding to increase the longevity of outgassed CO as the rate of outgassing would be unreasonably high to explain the masses observed if the CO were destroyed on short timescales (Kóspál et al. 2013; Kral et al. 2019).
@@
If the gas is primordial, the abundant H2 inherited from the protoplanetary disk could provide shielding (Kóspál et al. 2013). If the gas is secondary, CO can be replenished by outgassing from exocomets in the belt (Zuckerman and Song 2012), and shielding could be provided by CO itself (self-shielding) and/or atomic carbon (Matrà et al. 2017a; Kral et al. 2019), which could be produced by photodissociation of carbon-bearing parent molecules such as CO but also yet unseen species such as CO2 and CH4. Comparing the CO photodissociation timescale to the system age can be used to determine the gas origin scenario in CO bearing debris disks. Timescales longer than the system age allow for a primordial origin and shorter implies secondary gas.
@@
In CO-rich debris disks, assuming a $\frac{\text{CO}}{\text{H}_{2}}$ abundance ratio typical of protoplanetary disks leads to sufficient H2 to shield CO for several Myr, close to the system age (e.g. Kóspál et al. 2013). However, the same assumption for CO-poor belts leads to insufficient H2 to shield CO over the system age (e.g Marino et al. 2016; Matrà et al. 2017b; Marino et al. 2018; Matrà et al. 2019; Kral et al. 2019). This implies that gas in CO-poor systems, from younger systems like $\beta$ Pictoris (Matrà et al. 2017a) to older systems like Fomalhaut (Matrà et al. 2017b) and $\eta$ Corvi (Marino et al. 2016), is most likely secondary. This was corroborated by upper limits on the H2 density in the $\beta$ Pictoris disk obtained from non-local thermodynamic equilibrium (non-LTE) modelling of the CO excitation (Matrà et al. 2017a).
@@
In CO-rich belts, on the other hand, the origin of the gas remains an open question. These disks have CO masses comparable with the lower end observed in Herbig-Ae protoplanetary disks but around stars that are several Myr to a few 10s of Myr old (Moór et al. 2017). If the gas in these CO-rich disks is primordial, we would expect a large amount of H2, the most abundant molecular species in primordial gas. In protoplanetary disks, the gas phase CO-to-H2 ratio appears to decrease over time, producing $\frac{\text{CO}}{\text{H}_{2}}$ ratios between $10^{-4}$ and $10^{-6}$ (see Figure 5, Table 5 and Bergin and Williams (2017); Zhang et al. (2020)). This is because CO abundances start close to ISM-like, but over time CO can be processed into other carbon-carriers (such as CO2 and CH3OH) and at the same time CO can be sequestered onto icy grains as they grow and settle to the mid-plane (Schwarz et al. 2018; Bosman et al. 2018; Krijt et al. 2018, 2020).
On the other hand, for second generation gas, we may look to Solar System comets for comparison. Here, $\frac{\text{CO}}{\text{H}_{2}}$ ratios as high as 0.73 have been observed as H2 is mainly produced as a product of H2O photodissociation (Bockelée-Morvan and Biver 2017; Feldman et al. 2002). No molecules other than CO have thus far been observed in debris disks, thus it is unconfirmed whether exocometary ices should have compositions comparable to solar system cometary ice, but it is reasonable to expect a $\frac{\text{CO}}{\text{H}_{2}}$ abundance ratio higher than 10-4 in secondary gas.
@@
Some constraints on the presence of H2 in exocometary belts are present in the literature. For example, using HST Lecavelier des Etangs et al. (2001) constrained upper limits for the column density of H2 along the line of sight to the edge-on, CO-poor $\beta$ Pictoris disk to be ¡ 0.1 M and combined this with a CO column density (Roberge et al. 2000) to find a $\frac{\text{CO}}{\text{H}_{2}}$ ratio lower limit of $6\times 10^{-4}$, higher than expected of primordial gas. Their conclusion is in agreement with the conclusions of Matrà et al. (2017a) that a primordial origin scenario for the gas in $\beta$ Pictoris is unlikely as there is insufficient H2 present to provide shielding for the lifetime of the system. H2 can also be indirectly constrained by gas kinematics/distribution. Fernández et al. (2006) use the breaking experienced by ionised metallic gas due to unseen neutral H2 to conclude that the gas observed in $\beta$ Pictoris is likely secondary and produced in-situ due to evaporation of dust grains in the belt. HI Lyman $\alpha$ emission was detected by Wilson et al. (2017), but given that the hydrogen content of the disk is less than solar abundances, the authors conclude the HI likely originates from the dissociation of H2O released by cometary ices rather than H2 originating from the protoplanetary disk. Hughes et al. (2017) use the scale height of the gas in 49 Ceti to infer an upper limit for H2, via the mean molecular weight of the gas required by models to reproduce the observed scale height for CO. However, as pointed out by Marino et al. (2022) this is only valid if the CO scale height traces the scale height of the bulk of the gas, which may not be the case.
@@
Hot H2 emission has also been detected in systems bearing debris disks such as TWA 7 (Flagg et al. 2021) and AU Mic (Flagg et al. 2022), where the latter has no evidence for CO (Cronin-Coltsmann et al. 2023). It is unclear whether this hot H2 could originate from star spots or an inner disk of gas accreting onto the star, unrelated to the outer disks. Other works have attempted to detect warm $\text{H}_{2}$ with Spitzer, but these observations lacked the SNR/sensitivity to confirm or refute a primordial origin scenario (Chen et al. 2007; Kóspál et al. 2013).
@@
In this work, we aim to directly probe the column density of H2 within the CO-rich exocometary belts HD 110058 and HD 131488 for the first time and in doing so establish the CO origin. We expect any exocometary $\text{H}_{2}$ to be co-located with and similar temperature as the CO detected through UV absorption spectroscopy in both disks by Brennan et al. (2024) and in IR in this work. However it is worth noting that H2 should have a larger vertical extent than CO due to CO being more readily photodissociated at the disk’s surface and it’s lower mean molecular weight (Marino et al. 2020; Hughes et al. 2017). Both H2 and CO should be cold and produce narrow absorption lines from their ground vibrational energy levels. We focus on absorption spectroscopy in edge-on belts to use the star as a bright background continuum, and maximise the column density of gas along the line of sight to it. We identified the H2 v=1-0 S(0) line at 2223 nm as the most promising transition probing rovibrational absorption from the ground level of the H2 molecule. This is because, assuming the H2 gas in these systems has a similar temperature to the CO and is in LTE (predominantly collisionally excited), this line would be the strongest IR transition.
We used the newly upgraded, high-resolution CRIRES+ instrument on the VLT to observe the edge-on, CO-rich disks around HD 110058 and HD 131488. With a spectral resolution $R\sim 100000$ (Leibundgut et al. 2022), CRIRES+ is ideal for detecting line of sight absorption due to CO and H2. Our observations cover multiple v=2-0 rovibrational 12CO lines. Absorption from CO in edge-on debris disks has been successfully observed by Troutman et al. (2011), who reported low-$J$ CO absorption from the $\beta$ Pictoris disk using IRTF/CSHELL. Worthen et al. (2024) reported a non-detection of CO in absorption toward the edge-on HD 32297 disk using IRTF/iSHELL which was used to estimate the scale height of the disk and place an upper limit on the line-of-sight CO column density. The two disks in this study were measured in absorption in the UV using Hubble by Brennan et al. (2024). Our CRIRES+ observations allow for CO column density and temperature estimates for both HD 110058 and HD 131488.
@@
Stellar and disk properties for HD 110058 and HD 131488 are detailed in Table 1. Both are A-type stars (the most common spectral type for CO-rich belts, Moór et al. 2017) and thus provide bright stellar continua against which we searched for absorption. The disks in these systems are also very close to edge-on with inclinations of $85.5^{+2.5}_{-7.2}$ and $82\pm 3$ respectively (Moór et al. 2017; Hales et al. 2022). We chose these two young ($\sim 15$ Myr), CO-rich, edge-on belts as they have the highest CO column densities detected in absorption to date (Brennan et al. 2024).